# Nonlinear semiclassical spectroscopy of ultrafast molecular polariton dynamics

### v1

@@ Abstract
with the total signal simply being obtained as the sum of all contributions $\Delta T=\sum_{\vec{v}}\Delta T_{\vec{v}}$. One finds that there are two possible phase combinations contributing to $\alpha^{(2)(1)}$: $(0,1)$, and $(2,-1)$ [and the complex conjugate processes $(0,-1)$, $(-2,1)$], which are shown in Figs. 4(a), (b), and are associated with pump-probe and double-quantum coherence (DQC) spectroscopies, respectively. The first contribution [$(0,1)\equiv\pm\Phi_{p}\mp\Phi_{p}+\Phi_{p^{\prime}}$] is the ‘standard’ pump-probe, where populations and zeroth-order quantum coherences are created by twofold application of the pump with opposite phase [see Feynman diagrams in Fig. 4(c)]. The coherent superposition between upper and lower polaritons then undergo Rabi oscillations which are detected by the probe. The second phase combination [($(2,-1)\equiv 2\Phi_{p}-\Phi_{p^{\prime}}$] is negligible outside of the cavity, as it implies that the 2LS is excited twice. However, inside the cavity, since the pulses get extended on a time scale of $\kappa^{-1}$ [cf. Fig. 1(c)], the probe can act before the pump, therefore enabling this phase combination to arise within $\kappa^{-1}$. In the SM sm , we derive the phase contributions up to third order for an arbitrary multi-level system and provide an illustration for a slightly more complicated three-level system, where the $(2,-1)$-contribution (DQC) can also exist outside of the cavity.
@@
Discussion and conclusions.—We have presented an efficient and broadly applicable framework for nonlinear polariton spectroscopy, based on a mean-field evolution of the coupled light-matter system. The formalism extends conventional free-space spectroscopy by accounting for the feedback of the molecular polarization onto the cavity field and allows for an understanding of the polariton response beyond the stationary limit F. Ribeiro et al. (2018). Expanding the cavity field and molecular density matrix in terms of the input fields’ amplitudes and phases allows for a systematic understanding of the various contributions and excitation pathways to the nonlinear response.
@@
The formalism offers potential for extension and generalization via multiple avenues. For instance, while our current treatment relies on a semiclassical description, therefore restricting it to a coherent state of light within the cavity, a valuable extension could involve capturing nonclassical photon statistics, e.g., by using cumulant expansion techniques Fowler-Wright et al. (2023). Other areas of interest are the study of spatially-resolved nonlinear polariton spectroscopy, which would involve generalizing the zero-dimensional cavity Eq. (3) to the 3D set of cavity modes afforded by classical electromagnetic simulations Taflove and Hagness (2005); Yang et al. (2023); Balasubrahmaniyam et al. (2023); Xu et al. (2023), or cavities exhibiting inherent nonlinearities Finkelstein-Shapiro et al. (2023). Furthermore, while we have mainly focused on pump-probe spectroscopy involving two pulses, the formalism can also be readily applied to 2D or multidimensional coherent spectroscopy by considering three or more pulses driving the cavity. Finally, let us emphasize that our approach is not limited to molecular polaritons, but could also be extended to other strongly coupled light-matter platforms, such as exciton-polaritons in inorganic semiconductor microcavities Fryett et al. (2018) or atomic gases.

### v4

@@ Abstract
where $(0,-1)$ is the only phase contribution for the conjugate probe field sm and the total signal is obtained as $\Delta T=\sum_{\vec{v}}\Delta T_{\vec{v}}$. Two phase combinations contribute to $\alpha^{(2)(1)}$: $(0,1)$ and $(2,-1)$ [along with the complex conjugates $(0,-1)$ and $(-2,1)$], corresponding to pump–probe and double-quantum coherence (DQC) spectroscopies, respectively [Figs. 4(a,b)]. The pump–probe pathway $(0,1)\equiv\pm\Phi_{p}\!\mp\!\Phi_{p}\!+\!\Phi_{p^{\prime}}$ creates populations and zeroth-order quantum coherences via two pump pulses of opposite phase [Fig. 4(c)]. The DQC pathway $(2,-1)\equiv 2\Phi_{p}\!-\!\Phi_{p^{\prime}}$ is negligible outside the cavity because it requires the two-level system to be excited twice. However, inside the cavity, the probe can act on the pump-induced coherence within the lifetime of the bright state $(\kappa+\gamma_{\phi})^{-1}$ [left diagram in Fig. 4(d)], enabling this pathway to arise within this time interval. The right diagram in Fig. 4(d) does not contribute to the probe DT since the pump has not acted yet. Consequently, the $(2,-1)$ pathway isolates the bright state only, in contrast to the $(0,1)$ pathway, which accesses both bright and dark states (and probes only dephased dark states in the long time limit). Eq. (10) satisfies the phase matching conditions for degenerate 4WM with two pulses Yuen-Zhou et al. (2014).
@@
Discussion and conclusions.—We have presented an efficient and broadly applicable framework for nonlinear polariton spectroscopy, based on a mean-field evolution of the coupled light-matter system. The formalism extends conventional free-space spectroscopy by accounting for the feedback of the molecular polarization onto the cavity field and allows for an understanding of the polariton response beyond the stationary limit F. Ribeiro et al. (2018). Expanding the cavity field and molecular density matrix in terms of the input fields’ amplitudes and phases provides a systematic understanding of the various contribution to the nonlinear polariton response, and allows to separate purely bright state contributions from dark reservoir states.
@@
The formalism offers potential for extension and generalization via multiple avenues. For instance, while our current treatment relies on a semiclassical mean-field description, valuable extensions could involve capturing nonclassical photon statistics or many-body correlations among particles, e.g., by using cumulant or cluster expansion techniques Fowler-Wright et al. (2023); Axt and Stahl (1994). Other areas of interest are the study of spatially-resolved nonlinear polariton spectroscopy, which would involve generalizing the zero-dimensional cavity Eq. (3) to the 3D set of cavity modes afforded by classical electromagnetic simulations Gruetzmacher and Scherer (2003); Zhou et al. (2024); Yang et al. (2023); Balasubrahmaniyam et al. (2023); Xu et al. (2023), or cavities exhibiting inherent nonlinearities Finkelstein-Shapiro et al. (2023). Furthermore, while we have focused on pump-probe spectroscopy involving two pulses, the formalism can be readily applied to 2D or multidimensional coherent spectroscopy, which may also help to disentangle pure electronic dephasing from vibronic relaxation channels. Finally, let us emphasize that our approach is not limited to molecular polaritons, but could also be extended to other strongly coupled light-matter platforms, such as exciton-polaritons in inorganic semiconductor microcavities Fryett et al. (2018) or atomic ensembles.